Tohban Report 2014-03-19

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 12 Mar 2014
End Date: 19 Mar 2014
Tohban: Lindsay Glesener
Tohban email:
Next Tohban: Hazel Bain
List all reports


Solar Activity

Solar activity was a bit lower than the previous week. There are 10 ARs on the disk. AR 12010, in the SE quadrant, is currently the most active. In the past two days, activity has been picking up a little.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    32     3     0

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    23     2     0

And how many had EXCELLENT coverage?

 Flares above B, C, M, X class were      0     0     0     0

There were RHESSI flares/GOES flares 241 / 35 over the time range 12-Mar-14 19-Mar-14

Memory Management

No SSR issues this week. Max SSR fill was 26%, and it's been emptying with a pass or so to spare. We're still in Active / Vigorous. If activity stays as is we could probably stand to dial that back to Normal / Vigorous. However, since solar activity has mildly picked up in the last couple days, we'll hold off on this for now.

Spacecraft Status

Cold plate temperatures are 116.7K and 117.2K. Cryocooler is operating normally.

Spin rate is 14.5 rpm. It's time for a spinup, so this will be performed later today.

Data Gaps

No data gaps.

Detector issues

Thresholds were adjusted on several detectors. The front fast threshold was raised for Det 6, and lowered for Dets 4 and 2. Det 2 front slow threshold was also lowered.

Detector 2 seems to be operating ok with the lower thresholds. It is now completely segmented except for brief periods at the peak of the SAA.

A test was performed whereby detector 6's voltage was lowered for one orbit and then raised, though not all the way back up. Before beginning this test, G6 front fast valids were at 237k! During the test, with lowered HV, the rates dropped and livetime went up. When the HV was turned back up, rates increased again and livetime went down. It was decided to leave the detector at a bias voltage of 3063V, a net drop of 200V. Note that even at this value, fast rates were ~15k and livetime was ~77%. The minimum HV during the test was 2474V.

Based on the data following all this commanding, the decision was made the following day to raise the front fast threshold on D6 even further. G6 front fast was increased from 0xA0 to 0xB0 and G8 rear fast was increased from 0x70 to 0x80.

Command list from Monday, March 17:



14-076-20:34:17 /IDPUTABLE6 FRONTFASTDAC 14-076-20:34:33 /IDPULOAD VALUE=0xA0  ;Was 0x90


14-076-20:34:51 /IDPUTABLE4 FRONTFASTDAC 14-076-20:37:48 /IDPULOAD VALUE=22 (in decimal, commanded to 0x16) ;was 0x50 14-076-20:38:00 /IDPULOAD VALUE=0x22 (This was written to byte 9 (Rear slow)) ;was 0x30

14-076-20:38:07 /IDPUTABLE4 FRONTFASTDAC 14-076-20:38:12 /IDPULOAD VALUE=0x22 (correct front fast threshold change) 14-076-20:38:18 /IDPULOAD VALUE=0x30 (undoing the change to the rear slow)


14-076-20:38:58 /IDPUTABLE2 FRONTFASTDAC 14-076-20:39:22 /IDPULOAD VALUE=0x22 ;was 0x80



14-076-23:45:02 /IHVDAC DETECTOR=6, VOLTAGE=148 (lowered from 168) 14-076-23:46:05 /IHVDAC DETECTOR=6, VOLTAGE=128 (2474 V)


14-076-23:48:25 /IDPUTABLE2 FRONTSLOWDAC 14-076-23:48:40 /IDPULOAD VALUE=0x0C  ;was 0x38



14-077-01:23:14 /IHVDAC DETECTOR=6, VOLTAGE=138 (raised from 128) 14-077-01:24:40 /IHVDAC DETECTOR=6, VOLTAGE=148 14-077-01:26:08 /IHVDAC DETECTOR=6, VOLTAGE=158 (voltage ramp halted at 3064 V)

Command list from Tuesday, March 18:



14-077-23:30:17 /IDPUTABLE6 FRONTFASTDAC 14-077-23:30:46 /IDPULOAD VALUE=0xB0


14-077-23:31:31 /IDPUTABLE8 REARFASTDAC 14-077-23:34:26 /IDPULOAD VALUE=0x80

Other notes

Spacecraft Management

The yellow/red warnings and hard cutoffs for the cold plate two temperatures were all raised by 5 degrees.

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Disabled (4 minutes before and after eclipse)
Night time data (rears) Enabled
Require extra passes? No
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Detectors 6 and 9.
Personal tools